Positive definite functions on separable function spaces

Hrvoje Šikić, Zagreb, Croatia

Keywords: positive definite functions, semigroups, Laplace transforms, superprocesses AMS Subject classification: 43A35, 60J99

Abstract: A notion of a separable function space is defined. It is proved that a positive definite function on the semigroup of the positive elements of a separable function space is the Laplace transform of a measure concentrated on the set of positive functionals.